OpenStudy (lifeisadangerousgame):

Help ASAP would be awesome c:

OpenStudy (lifeisadangerousgame):

I need to solve x - 3y + 2z = 1 -x + 4y - 4z = 1 5x - y + 6z = 9

OpenStudy (tkhunny):

Favorite method?

OpenStudy (lifeisadangerousgame):

So far what I did is x - 3y + 2z = 1 +(-x + 4y - 4z = 1) and I got y - 2z = 1 So then I multiplied -x + 4y - 4z = 1 by 5 to get -5x + 20y - 20z = 5 Then I did this -5x + 20y - 20z = 5 + 5x - y + 6z = 9 to get 19y - 14z = 14 Am I correct so far? I don't have a favorite method

hartnn (hartnn):

1+1 = 2 isn't it ?

OpenStudy (lifeisadangerousgame):

ahh I missed that, yes it does ><

OpenStudy (lifeisadangerousgame):

that fails miserably I know >.>

hartnn (hartnn):

2nd equation is correct 19y-14z=14

OpenStudy (joannablackwelder):

Looks good to me. Now you just need to take the two variable equations and solve by substitution or elimination.

hartnn (hartnn):

y-2z = 2 so, y = 2+2z plug this in 2nd correct equation

OpenStudy (lifeisadangerousgame):

So what I'm doing next is multiplying y - 2z = 2 by -7 so that it's -7y + 14z = 14 so that I can eliminate "z" in this: 19y - 14z = 14 + -7y + 14z = -14 which gives me 12y = 0 Is that right?

hartnn (hartnn):

yes, y comes out to be 0 good :)

hartnn (hartnn):

now you can find z easily

hartnn (hartnn):

put y= 0 in y-2z=2

OpenStudy (lifeisadangerousgame):

I need to find x as well, can I find z first or do I have to find x and then y?

hartnn (hartnn):

you will be able to find z first, not x because you have one equation in y and z

hartnn (hartnn):

**two equations in y and z use any

OpenStudy (lifeisadangerousgame):

I'll put y into this equation "y-2z = 2" so y-2z = 2 0 - 2z = 2 -2z = 2 divide by -2 on both sides z = -1?

hartnn (hartnn):

yes, z=-1 is correct :) now find x using any equation in x

OpenStudy (lifeisadangerousgame):

I have to use an original equation this time correct?

hartnn (hartnn):

yes, any 1 of the 3 original equations...

OpenStudy (lifeisadangerousgame):

Okay I'll use "5x - y + 6z = 9" So 5x - y + 6z = 9 5x - 0 + 6(-1) = 9 5x - 6 = 9 +6 +6 5x = 15 divide by 5 on both sides x = 3

hartnn (hartnn):

correct! :)

OpenStudy (lifeisadangerousgame):

Awesome! Thank you! :3